Let N be a Poisson random variable with parameter λ. Suppose ξ1,ξ2,... is a sequence of i.i.d. random variables with mean µ and variance σ2, independent of N. Let SN = ξ1 + ...ξN. Determine the mean and variance of SN
Answer #1

The expected value of a compound Poisson process can be calculated using a result known as Wald's equation as:

{displaystyle ,E(Y(t))=E(D_{1}+...+D_{N(t)})=E(N(t))E(D_{1})=E(N(t))E(D)=lambda tE(D).}

Making similar use of the law of total variance, the variance can be calculated as:

var(Y(t ))-E(var(y(t) V(t))) + var(E(y(t) V(t))) E(N(t) var(D)) +var(N(t)E(D var(D)E(NtE(D)2 var(N(t)) At(var(D) + E(D)2)

E(Sn) = λ * µ

Var(Sn) = λ * (µ ^2 + σ^2)